Write the slope-intercept form of the equation of the line through (-1,2) that is perpendicular to the line x+3y=4 hope you can help, thanks
Answer by checkley71(8403)

FIND THE SLOPE OF THE EQUATION X+3Y=4 THUS
3Y=-X+4
Y=(-1/3)X+4/3 THUS THE SLOPE IS -1/3
A LINE PERPENDICULAR TO THIS LINE MUST HAVE A SLOPE THAT IS THE NEGATIVE RECIPRICLE OF THIS SLOPE THUS THE NEW SLOPE IS 3
SO Y=3X+b NOW SUBSTITUTING THE X & Y VALUS FOR THE POINT WE GET THE Y INTERCEPT
2=3*-1+b
2=-3+b
2+3=b
5=b THUS WE HAVE THE FORMULAR FOR THIS LINE
Y=3X+5 THAT HAS A SLOPE OF 3, A Y INTERCEPT OF 5 & PASSES THROUGH THE POINTS
(-1,2)